(3x^2+2xy+y^3)dx + (x^2+y^2)dy=0
1
Expert's answer
2020-11-09T20:16:07-0500

"\begin{aligned}
(3x^2y + 2xy + y^3)dx &+ (x^2 + y^2) dy = 0\\
\\
M=3x^2y + 2xy + y^3 &; \quad N = x^2 + y^2\\
M_y = 3x^2 + 2x + 3y^2 &; \quad N_x = 2x\\
\\
\therefore \dfrac{M_y-N_x}{N} = 3\\
\\
Thus,\ \mu(x)= e^{\int{3dx}} &= e^{3x}

\end{aligned}"Employing "\mu(x)" into the general equation, we have;

"\begin{aligned}
e^{3x}(3x^2y + 2xy + y^3)dx &+ e^{3x}(x^2 + y^2) dy = 0\\


\end{aligned}"which, by construction, must be exact. So we seek a function "\psi" such that;

"\\frac{\\partial\\psi}{\\partial x} =e\r^{3x}(3x\r^2y + 2xy + y\r^3\r)\\\\\n\\frac{\\partial \\psi\r}{\\partial y} = e\r^{3x}\r(x\r^2 + y\r^2\r)"


Integrating the first equation with respect to x and the second equation with respect to y, we get;

"\psi (x, y) = x^2ye^{3x} + \frac{1}{3}y^3e^{3x}+h_1(y)\\
\psi (x, y) = x^2ye^{3x} +\frac{1}{3}y^3e^{3x}+h_2(x)"


Comparing the two expressions above we observe that we must take a "h_1y = h_2x =" C (a constant).

Thus, function "\psi" satisfying the exact function must be of the form;


"\\psi (x, y) = e\r^{3x}x^2y + e\r^{3x}y^3 + C"


Thus, the general solution to the equation is of the form;

"e\r^{3x}x\r^2y + e\r^{3x\r}y^3 = C"